# Thermal circuit model for silicon quantum-dot array structures

### v1

@@ III.1 Device structure
To verify the proposed thermal model, we experimentally measure $T_{e}$ in a QD array. Figure 3(a) shows the device structure Utsugi et al. (2023). In this device, we can measure the temperature at different distances from the heating point. The T-shaped silicon channel (green) is fabricated using silicon on insulator (SOI), and multiple polysilicon gate electrodes are formed on the T-shaped Si channel. The horizontal channel is covered with first gates (BG0-BG3, light blue) and second gates (PG0-PG3, blue), for a total of seven gates. These fine-pitch gate structures are fabricated using the self-align patterning process Lee et al. (2020); kun . Each gate has terminals on both sides, allowing current to flow through them by applying DC voltages through wiring connected to instruments at room temperature.
@@
The four first gates (BG0-BG3) are used as heat sources by flowing current through each gate. Figure 3(b) shows the cross-section between A and A’ in Fig. 3(a). Corresponding to the model shown in Fig. 2(a), the unit cell length of this device is $L_{\rm{cell}}=120$ nm (we also measure the device of $L_{\rm{cell}}=160$ nm, see Appendix B).
The silicon chip with the above structure is glued to a cryogenic printed circuit board (QBoard, Qdevil) using silver paste, and each terminal is wired using aluminum bonding wire. Measurements were performed using a dilution refrigerator (Proteox, Oxford Instruments) with a base temperature of 8 mK.
@@ III.2 Measurement
We use CBT to estimate the local temperature at the SET. In CBT, there are several restrictions for QD parameters for accurately measuring temperature, e.g., the tunnel rate of barriers of SET $h\Gamma$, energy spacing of the quantum levels $\Delta E$, thermal energy $k_{B}T$, and charging energy $e^{2}/C$  Beenakker (1991). In this experiment, we use the classical regime, i.e., $h\Gamma\ll\Delta E\ll k_{B}T\ll e^{2}/C$. By measuring the current $I_{d}$ flowing through the SET while sweeping the voltage of the SGS gate $V_{g}$ and the drain $V_{d}$ [Fig. 4(a)], we can obtain the SGS gate voltage dependence of the differential conductance $G=dI_{d}/dV_{d}$ [Fig. 4(b)] and estimate the electron temperature around SET $T_{e}$. The differential conductance $G$ in the region where $V_{d}$ is small is Beenakker (1991)
@@
where $G_{\rm{max}}$, $\alpha$, $e$, $k_{B}$, and $V_{0}$ are the peak value of $G$, the lever arm, the elementary charge, the Boltzmann constant, and the voltage value at $G_{\rm{max}}$, respectively. The lever arm was estimated as $\alpha=0.08$ by obtaining the Coulomb diamond shown in Fig. 4(a). Figure 4(c) and 4(d) show examples of $G$ (Coulomb peak) at different $T_{e}$ situations. It can be confirmed that the Coulomb peak shape reflects the change in $T_{e}$, and $T_{e}$ can be estimated by fitting the experimental result by the curve of Eq. (8). This measurement method is constrained by the tunnel rate and can be used in the range $k_{B}T_{e}\gg h\Gamma$, where $h\Gamma/k_{B}\sim 1$ K in this experiment.
@@
We use a local heater by applying a current to each of the four gate electrodes BG0-BG3 as shown in Fig. 5(a). If the electric field around the SET changes during the measurement, the SET conditions are affected by the electric field and the temperature measurement cannot be performed correctly. Therefore, we apply a voltage of $+V/2$ and $-V/2$, where $V=\sqrt{RP}$ ($R$ is the resistance of the entire gate wiring) to each terminal of the gate to allow current to flow through the gate. Due to the symmetry of the wiring structure in the dilution refrigerator, we set the voltage to be approximately zero on the channel. This stabilizes the SET condition because the voltage condition around the SET is stable even if the heating power $P$ is changed by varying $V$, where $P=V^{2}/R$. We measure the resistance of each gate wiring excluding the additional resistance on the circuit board for a low pass filter and use these values to calculate $P$, considering that the polysilicon gate has the dominant resistance among all of the gate wiring.
Figure 4: (a,b) Coulomb diamond of (a) the drain current of SET $I_{d}$ and (b) the differential conductance of SET $G$. The (a) white and (b) black dotted lines show one of the Coulomb diamond areas. (c) and (d) show examples of Coulomb peak and fitting curve with Eq. (8) at different temperatures. The estimated temperatures are (c) 2.4 K and (d) 5.5 K.
Figure 5(b) shows the experimental results of $T_{e}$ as a function of the heating power $P$ when each gate is used as a heat source. As expected, $T_{e}$ increases monotonically with $P$ and $1/D$. The experimental results can be fitted using Eq. (7) as
$T_{e}=A_{i}\sqrt{P+P_{0}},$ (9)
@@
We measure another device of different gate pitch as shown in Appendix B. From these results, we conclude that the proposed model successfully reproduces the experimental results. This demonstrates its capability to model the thermal conduction characteristics of silicon QD array structures effectively.
Figure 5: (a) Schematic of local heaters by flowing current in barrier gates. The resistance of each gate excluding the resistance of the low pass filter in Qboard at cryogenic temperatures are $R_{\rm{BG0}}=78.8~{}\rm{k}\Omega$, $R_{\rm{BG1}}=78.1~{}\rm{k}\Omega$, $R_{\rm{BG2}}=70.9~{}\rm{k}\Omega$ and $R_{\rm{BG3}}=61.3~{}\rm{k}\Omega$, respectively. (b) Experimental results of $T_{e}$ as a function of $P$ and $D$. The dotted curves are the best-fit results with Eq. (9). The inset shows the best-fit results of $A$ as a function of $D$ and the dotted curve is the best-fit result with $A=ae^{-\frac{\gamma_{\rm{th}}}{2}D}$, where $a=8.5~{}\rm{K}/\sqrt{\rm{\mu W}}$ and $2/\gamma_{\rm{th}}=280~{}\rm{nm}$.
@@ IV Discussion
To consider the limitations of our model, we re-summarize the assumptions for deriving Eq. (4) according to the above discussions: (i) The thermal circuit of the model is an infinite periodic structure, (ii) $R_{\rm{in}}$ and $R_{\rm{ex}}$ are temperature dependent and spatially uniform, and (iii) the net effective temperature of the QD array is proportional to $T_{e}$. Moreover, for deriving Eq. (5), we also assume that the temperature dependence of $R_{\rm{in}}$ and $R_{\rm{ex}}$ are metallic ($\beta_{\rm{in}}\approx\beta_{\rm{ex}}\approx 1$). Despite these potential sources of modeling error, the consistency between the experimental results and the model supports the validity of our simplifications. The model validity is further investigated using the order estimation in Appendix C.
@@
We also discuss how to design the QD array structure. To mitigate the thermal effects on the qubits, based on Eq. (5), the following conditions should be met: (I) $P$ should be reduced, (II) $D$ should be increased, and (III) the device-dependent parameters $a$ and $2/\gamma_{\rm{th}}$ should be minimized. Conditions (I) and (II) are straightforward, i.e., the amount of heating power needs to be reduced, for example, by lowering the electrical resistance of the gate, and the qubits need to be kept appropriately distant from the heat sources. Our main result is condition (III), which is useful for quantitative thermal design of the QD array structures. Reducing $a$ implies that even if the heating power is high, the temperature rise is low; in other words, the structure does not easily accumulate heat. To reduce $a\propto\sqrt[4]{R_{\rm{in}}R_{\rm{ex}}}$, the thermal resistance of the entire structure should be minimized. On the other hand, reducing $2/\gamma_{\rm{th}}$ (the characteristic length) means creating a structure that makes it difficult to transport heat to the qubits. To reduce $2/\gamma_{\rm{th}}\approx 2\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$, the ratio of heat dissipation to heat inflow should be increased. For example, an additional metallic heat sink at the top and bottom of the gate layer can be effective. This can be achieved using technologies such as through-silicon vias (TSVs) Taguchi et al. (2024).
@@
In future work, we will investigate the thermal dynamic model of the QD array structure to control the heat sources using real-time ambient temperature measurements. The thermal dynamic model is also strongly related to the operation fidelity of qubits, including the heat-induced frequency shift Takeda et al. (2018); Undseth et al. (2023); Sato and Kawahara (2024). This will be modeled by the heat capacitance $C_{\rm{ex}}$ and measured by a high-speed readout method such as reflectometry Champain et al. (2024).
@@ Appendix A Thermophysical properties
We summarize thermophysical properties of represented materials in the typical silicon fabrication process based on Refs. Duthil (2014); Ekin (2006); ais , as shown in Fig. 6.
@@
Appendix B Measurement result of device B
We measure another device with $L_{\rm{cell}}=160$ nm (device B), similar to the device described in the main text (device A). Figure 7 shows the results, which exhibit the same trend as those in Fig. 5. However, the electron temperature is lower than that of device A. This is attributed to the larger gate pitch, which implies a longer gate length and a higher ratio of the polysilicon gate to the insulator. This configuration likely decreases the resistance $R_{\rm{in}}$ while maintaining $R_{\rm{ex}}$, resulting in an increased thermal characteristic length $2/\gamma_{\rm{th}}\approx 2\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$. On the other hand, $a\propto\sqrt[4]{R_{\rm{in}}R_{\rm{ex}}}$ changes minimally due to the fourth root. This measurement further validates the proposed thermal model. In Appendix C, we further summarize the experimental results.
Figure 7: Experimental results of $T_{e}$ as a function of $P$ and $D$ in device B of $L_{\rm{cell}}=160$ nm. The dotted curves are the best-fit results with Eq. (9). The inset shows the best-fit results of $A$ as a function of $D$ and the dotted curve is the best-fit result with $A=ae^{-\frac{\gamma_{\rm{th}}}{2}D}$, where $a=7.0~{}\rm{K}/\sqrt{\rm{\mu W}}$ and $2/\gamma_{\rm{th}}=700~{}\rm{nm}$.
Appendix C Summary of experimental results and order estimation of heat inflow resistance
Device A Device B
$L_{\rm{cell}}$ (nm) 120 160
$a$ (K$/\sqrt{\rm{\mu W}}$) 8.5 7.0
$2/\gamma_{\rm{th}}$ (nm) 280 700
$R_{\rm{in}}T_{e}$ ($\rm{K^{2}/(m\cdot W)}$) $5.1\times 10^{14}$ $1.4\times 10^{14}$
$R_{\rm{ex}}T_{e}$ ($\rm{K^{2}\cdot m/W}$) $10$ $17$
Table 1: The device-dependent parameters calculated from the experimental results, and order estimation results (the last row).
Table 1 presents the device-dependent parameters estimated from the experiments, where we compare devices A (main text) and B (Appendix B). To evaluate the thermal resistances, we use the relations of $R_{\rm{in}}T_{e}=a^{2}\gamma_{\rm{th}}$ and $R_{\rm{ex}}T_{e}=a^{2}/\gamma_{\rm{th}}$ derived from Eqs. (2) and (3), where $\beta=1$ is used. We analyze these results to the order estimation values of heat inflow resistance $R^{\rm{est}}_{\rm{in}}$ on the basis of the device structures and the thermophysical properties shown in Fig. 6 (see Appendix B). Note that the order estimation of $R_{\rm{ex}}$ is difficult due to the complicated heat dissipation path.
Using the model of the QD array structure shown in Fig. 2 analogous to the series resistance, we estimate $R^{\rm{est}}_{\rm{in}}$ as
@@ Appendix C Summary of experimental results and order estimation of heat inflow resistance
where $S$ is the cross-sectional area of the heat inflow path, $L_{\rm{PolySi}}$ ($L_{\rm{SiO_{2}}}$) is the total length of the polysilicon gates BG and PG (SiO2), where $L_{\rm{cell}}=L_{\rm{PolySi}}+L_{\rm{SiO_{2}}}$, and $\kappa_{\rm{PolySi}}$ ($\kappa_{\rm{SiO_{2}}}$) is the thermal conductivity of the polysilicon gates (SiO2). We use $S=10^{-13}$ $\rm{m}^{2}$, where $S$ is calculated as the product of the width ($W$) and height ($H$) of the heat inflow path and we use $W=1~{}\rm{\mu m}$ and $H=0.1~{}\rm{\mu m}$. From the device structure, we use $L_{\rm{PolySi}}=90$ nm for device A and $L_{\rm{PolySi}}=130$ nm for device B, while $L_{\rm{SiO_{2}}}=30$ nm for both devices. From the thermophysical properties shown in Fig. 6, we use $\kappa_{\rm{PolySi}}=0.05~{}\rm{W/(m\cdot K)}$ and $\kappa_{\rm{SiO_{2}}}=0.01~{}\rm{W/(m\cdot K)}$ assuming around 1 K. We obtain the estimation results as $R^{\rm{est}}_{\rm{in}}=4.0\times 10^{14}$ for device A and $R^{\rm{est}}_{\rm{in}}=3.5\times 10^{14}$ for device B at $T_{e}=1$ K, as shown in Table 1. The lower $R_{\rm{in}}$ for device B than for device A is consistent with the experimental trend. This result demonstrates the validity of our model. Additionally, the order of $R^{\rm{est}}_{\rm{in}}$ values are comparable to the experimental results at $T_{e}=1$ K, validating the formulation in Eq. 10 at least for the order estimation. We also expect that by improving this simple model or combining it with different approaches, e.g., finite element simulations, we can develop a framework for predicting thermal characteristics more reliably and accurately.

### v2

@@ III.1 Device structure
To verify the proposed thermal model, we experimentally measure $T_{e}$ in a QD array. Figure 7(a) shows the device structure Utsugi et al. (2023). In this device, we can measure the temperature at different distances from the heating point. The T-shaped silicon-on-insulator (SOI) channel (green) is fabricated, and multiple polysilicon gate electrodes are formed on the T-shaped Si channel. The horizontal channel is covered with first gates (BG0-BG3, light blue) and second gates (PG0-PG3, blue), for a total of seven gates. These fine-pitch gate structures are fabricated using the self-align patterning process Lee et al. (2020); Kuno et al. (2025). Each gate has terminals on both sides, allowing current to flow through them by applying DC voltages through wiring connected to instruments at room temperature.
@@
The four first gates (BG0-BG3) are used as heat sources by flowing current through each gate. Figure 7(b) shows the cross-section between A and A’ in Fig. 7(a). Corresponding to the model shown in Fig. 2(a), the unit cell length of this device is $L_{\rm{cell}}=120$ nm (we also measure the device of $L_{\rm{cell}}=160$ nm).
The silicon chip with the above structure is glued to a cryogenic printed circuit board (QBoard, Qdevil) using silver paste, and each terminal is wired using aluminum bonding wire. Measurements are performed using a dilution refrigerator (Proteox, Oxford Instruments) with a base temperature of 8 mK.
@@ III.2 Measurement
[…]
By measuring the current $I_{d}$ flowing through the SET while sweeping the voltage of the SGS gate $V_{g}$ and the drain $V_{d}$ [Fig. 4(a)], we can obtain the SGS gate voltage dependence of the differential conductance $G=dI_{d}/dV_{d}$ [Fig. 4(b)] and estimate the electron temperature around SET as $T_{e}$. Note that the Coulomb diamonds shown in Fig. 4(a) and 4(b) deviate from the ideal shape, indicating the possible formation of multiple QDs near or within the SET. This is likely due to structural imperfections in the device, such as disorder or surface roughness in the SET channel. To mitigate the influence of such imperfections, we select a Coulomb diamond that is as well-isolated as possible. Furthermore, the bias offset in the source-drain voltage $V_{\rm{ds}}$ observed in Fig. 4(a) and 4(b) originates from imperfections in the measurement system—specifically, an offset in the current measurement. This may be attributed to the Seebeck effect in the cables or fluctuations in the ground level in the source measure unit. However, this offset does not significantly affect the measurement results, as it is appropriately corrected by $V_{\rm{correct}}$ during data analysis, i.e., $V_{d}=V_{\rm{sd}}-V_{\rm{correct}}$.
The differential conductance $G$ in the region where $V_{d}$ is small is Beenakker (1991)
@@
where $G_{\rm{max}}$, $\alpha$, $e$, $k_{B}$, and $V_{0}$ are the peak value of $G$, the lever arm, the elementary charge, the Boltzmann constant, and the voltage value at $G_{\rm{max}}$, respectively. The lever arm was estimated as $\alpha=0.08$ by obtaining the Coulomb diamond shown in Fig. 4(a). Figure 4(c) and 4(d) show examples of $G$ (Coulomb peak) at different $T_{e}$ situations. It can be confirmed that the Coulomb peak shape reflects the change in $T_{e}$, and $T_{e}$ can be estimated by fitting the experimental result by the curve of Eq. (8). This measurement method is constrained by the tunnel rate and is applicable in the regime where $k_{B}T_{e}\gg h\Gamma$. Given that $h\Gamma/k_{B}<0.3~\mathrm{K}$ in this experiment, the condition $k_{B}T_{e}\gg h\Gamma$ implies that the method is applicable for $T_{e}\gtrsim 1~\mathrm{K}$.
@@
We use a local heater by applying a current to each of the four gate electrodes BG0-BG3 as shown in Fig. 5(a). If the electric field around the SET changes during the measurement, the SET conditions are affected by the electric field and the temperature measurement cannot be performed correctly. Therefore, we apply a voltage of +V/2+V/2 and V/2-V/2, where $V=\sqrt{RP_{0}}$ ($R$ is the resistance of the entire gate wiring) to each terminal of the gate to allow current to flow through the gate. Due to the symmetry of the wiring structure in the dilution refrigerator, we set the voltage to be approximately zero on the channel. This stabilizes the SET condition because the voltage condition around the SET is stable even if the heating power $P_{0}$ is changed by varying $V$, where $P_{0}=V^{2}/R$. We measure the resistance of each gate wiring excluding the additional resistance on the circuit board for a low pass filter and use these values to calculate $P_{0}$, considering that the polysilicon gate has the dominant resistance among all of the gate wiring.
Figure 4: (a,b) Coulomb diamond of (a) the drain current of SET $I_{d}$ and (b) the differential conductance of SET $G$. The (a) white and (b) black dotted lines show one of the Coulomb diamond areas. (c) and (d) show examples of Coulomb peak and fitting curve with Eq. (8) at different temperatures. The estimated temperatures are (c) 2.4 $\pm~0.1$ K and (d) 5.5 $\pm~0.1$ K, where the estimated error is the standard deviation of the $T_{e}$ estimate. It should be noted that the increase in the conductance peak value in (c) and (d) is most likely attributed to temperature-induced changes in the tunneling rate of the SET formed by the multiple QDs. The bias offset of source-drain voltage $V_{\rm{ds}}$ in (a) and (b) originates from imperfections in the measurement system, as described in the main text. This offset is appropriately corrected in the estimation of $T_{e}$.
Figure 5(b) shows the experimental results of $T_{e}$ as a function of the heating power denoted by $P_{0}$ and the distance between the heater and the SET denoted by $D$ when each gate is used as a heat source. As expected, $T_{e}$ increases monotonically with $P_{0}$ and $1/D$. The experimental results can be fitted using Eq. (5) as
$T_{e}=A_{i}(P_{0}+P_{B})^{\frac{1}{\beta+1}},$} (9)
where $(i\in\{$BG0,BG1,BG2,BG3})\}) [see the dotted curves in Fig. 5(b)]. The inset of Fig. 5(b) plots $A_{i}$ and shows the best-fit curve with $A_{i}=ae^{-\frac{D}{L_{\rm{th}}}}$ from Eq. (5). We experimentally confirm that $T_{e}$ rises is proportional to $P_{0}^{\frac{1}{\beta+1}}$ as in the model, and the heating effect of distance is reduced exponentially as expected. However, we note that it is difficult to determine from this experiment whether the fitted curve strictly follows exponential decay. To clarify this point, it would be necessary to either increase the number of measurement conditions or use a setup in which the value of $\beta$ is known—i.e., independently measured in advance by an appropriate method. We also measure another device of different gate pitch as shown in Fig. 5(c). From these results, we conclude that the proposed model successfully reproduces the experimental results. This demonstrates its capability to model the thermal conduction characteristics of silicon QD array structures effectively.
@@ IV Discussion
To consider the limitations of our model, we re-summarize the assumptions for deriving Eq. (5) according to the above discussions: (i) The thermal circuit of the model is a semi-infinite periodic structure. (ii) Both $R_{\rm{in}}(x)$ and $R_{\rm{ex}}(x)$ are temperature-dependent and spatially non-uniform, (iii) The local effective temperature of the QD array is proportional to $T_{e}(x)$. (iv) $T_{e}(x)\gg T_{\rm{base}}$. (v) $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\,(\equiv\beta)$. Regarding assumption (i), numerical simulations show that even a finite and discrete circuit exhibits the same dependence as Eq. (5) (see Supplementary Material III). Despite these potential sources of modeling error, the consistency between the experimental results and the model supports the validity of our simplifications.
Device A Device B
$L_{\rm{cell}}$ (nm) 120 160
[…]
$L_{\rm{th}}$ (nm) 264 500
$R_{\rm{in}}T_{e}$ ($\rm{K^{2}/(m\cdot W)}$) $2.7\times 10^{14}$ $1.2\times 10^{14}$
$R_{\rm{ex}}T_{e}$ ($\rm{K^{2}\cdot m/W}$) $9.5$ $15.6$
Table 1: The device-dependent parameters calculated from the experimental results assuming $\beta=1$, and order estimation results at $T_{e}=1$ K (the last row).
Table 1 presents the device-dependent parameters estimated from the experiments, where we compare devices A and B. To evaluate the thermal resistances, we use the relations of $R_{\rm{in}}T_{e}^{\beta}=a^{\beta+1}/L_{\rm{th}}$ and $R_{\rm{ex}}T_{e}^{\beta}=a^{\beta+1}L_{\rm{th}}/(\beta+1)$ derived from the definitions of the $a$ and $L_{\rm{th}}$. Here we use $\beta=1$. We analyze these results to the order estimation values of heat inflow resistance $R^{\rm{est}}_{\rm{in}}$ on the basis of the device structures and the thermophysical properties (see Supplementary Material II). We note that the order estimation of $R_{\rm{ex}}$ is difficult due to the complicated heat dissipation path. We have successfully quantified the thermal characteristics of two QD array structures with different gate pitches. The results indicate that the two devices differ significantly in the characteristic length $L_{\rm{th}}$, while showing only minor variation in the parameter $a$. We believe that further experimental investigations to clarify the origin of this discrepancy will contribute to the optimization of the device structure.
@@
We also discuss how to design the QD array structure. To mitigate the thermal effects on the qubits, based on Eq. (5), the following conditions should be met: (I) $P_{0}$ should be reduced, (II) $D$ should be increased, and (III) the device-dependent parameters $a$ and $L_{\rm{th}}$ should be minimized. Conditions (I) and (II) are straightforward, i.e., the amount of heating power needs to be reduced, for example, by lowering the electrical resistance of the gate, and the qubits need to be kept appropriately distant from the heat sources. Our main result is condition (III), which is useful for quantitative thermal design of the QD array structures. Reducing $a$ implies that even if the heating power is high, the temperature rise is low; in other words, the structure does not easily accumulate heat. To reduce $a\propto\sqrt[2(\beta+1)]{R_{\rm{in}}R_{\rm{ex}}}$, the thermal resistance of the entire structure should be minimized. On the other hand, reducing $L_{\rm{th}}$ means creating a structure that makes it difficult to transport heat to the qubits —i.e., thermally separating the heat source from the qubits. To reduce $L_{\rm{th}}\propto\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$, the ratio of heat dissipation to heat inflow should be increased. For example, an additional metallic heat sink at the top and bottom of the gate layer can be effective. This can be achieved using technologies such as through-silicon vias (TSVs) Taguchi et al. (2024).
@@
In future work, we will investigate the thermal dynamic model of the QD array structure to control the heat sources using real-time ambient temperature measurements. The thermal dynamic model is also strongly related to the operation fidelity of qubits, including the heat-induced frequency shift Takeda et al. (2018); Undseth et al. (2023); Sato and Kawahara (2024). This will be modeled by introducing the heat capacitance in the thermal circuit and measured by a high-speed readout method such as reflectometry Champain et al. (2024). The simplest case of such a dynamic thermal effect is the heating caused by gate pulsing. In this case, the introduction of heat capacitance leads to a thermal time constant, defined as the product of local thermal resistance and local heat capacitance on the thermal circuit, which governs the transient thermal response. If the time constant is shorter than the pulse duration, the thermal behavior is expected to resemble the static case, and the thermal coupling can still be characterized by $L_{\rm{th}}$. Conversely, if the time constant is longer than the pulse duration, the thermal impact is expected to be reduced. This behavior is analogous to a low-pass filter response in electrical circuits.
Another compelling topic is phonon engineering Maire et al. (2017). If we can design a device structure in which phonons propagate coherently as waves, it may be possible to achieve more sophisticated thermal control. We intend to further investigate this concept and consider how our proposed model could be situated within the broader context of phonon engineering.
@@
Supplementary Material
The Supplementary Material provides additional information to support the discussion in the main text.
@@ Acknowledgements.
AUTHOR DECLARATIONS
Conflict of interest
The authors have no conflicts to disclose.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Thermophysical properties
We summarize thermophysical properties of represented materials in the typical silicon fabrication process based on Refs. Duthil (2014); Ekin (2006); ais , as shown in Supplementary Fig. 6.
@@
Appendix B Order estimation of heat inflow resistance
Using the model of the QD array structure shown in Fig. 2 in the main text analogous to the series resistance, we estimate $R^{\rm{est}}_{\rm{in}}$ as
@@ Appendix B Order estimation of heat inflow resistance
where $S$ is the cross-sectional area of the heat inflow path, $L_{\rm{PolySi}}$ ($L_{\rm{SiO_{2}}}$) is the total length of the polysilicon gates BG and PG (SiO2), where $L_{\rm{cell}}=L_{\rm{PolySi}}+L_{\rm{SiO_{2}}}$, and $\kappa_{\rm{PolySi}}$ ($\kappa_{\rm{SiO_{2}}}$) is the thermal conductivity of the polysilicon gates (SiO2). We use $S=10^{-13}$ $\rm{m}^{2}$, where $S$ is calculated as the product of the width ($W$) and height ($H$) of the heat inflow path and we use $W=1~\rm{\mu m}$ and $H=0.1~\rm{\mu m}$. From the device structure, we use $L_{\rm{PolySi}}=90$ nm for device A and $L_{\rm{PolySi}}=130$ nm for device B, while $L_{\rm{SiO_{2}}}=30$ nm for both devices. From the thermophysical properties shown in Fig. 6, we use $\kappa_{\rm{PolySi}}=0.05~\rm{W/(m\cdot K)}$ and $\kappa_{\rm{SiO_{2}}}=0.01~\rm{W/(m\cdot K)}$ assuming around 1 K. We obtain the estimation results as $R^{\rm{est}}_{\rm{in}}=4.0\times 10^{14}$ for device A and $R^{\rm{est}}_{\rm{in}}=3.5\times 10^{14}$ for device B at $T_{e}=1$ K, as shown in Table 1 in the main text. The lower $R_{\rm{in}}$ for device B than for device A is consistent with the experimental trend. This result demonstrates the validity of our model. Additionally, the order of $R^{\rm{est}}_{\rm{in}}$ values are comparable to the experimental results at $T_{e}=1$ K, validating the formulation in Eq. (10) at least for the order estimation. We also expect that by improving this simple model or combining it with different approaches, e.g., finite element simulations, we can develop a framework for predicting thermal characteristics more reliably and accurately.
Appendix C Finite and descrete thermal circuit model
We present a finite and discrete thermal circuit model, analyzed using the ABCD matrix method. The ABCD matrix (also known as the transmission matrix) describes the relationship between input and output temperatures (analogous to voltages) and powers (analogous to currents) in a series of two-port networks shown in Supplementary Fig. 7(a) as
$\begin{bmatrix}T_{e}[0]\\ P[0]\end{bmatrix}$ $=\begin{bmatrix}1+\frac{R_{\rm{in}}[1]}{R_{\rm{ex}}[1]}&R_{\rm{in}}[1]\\ \frac{1}{R_{\rm{ex}}[1]}&1\end{bmatrix}\cdots\begin{bmatrix}1+\frac{R_{\rm{in}}[n]}{R_{\rm{ex}}[n]}&R_{\rm{in}}[n]\\ \frac{1}{R_{\rm{ex}}[n]}&1\end{bmatrix}\begin{bmatrix}T_{e}[n]\\ P[n]\end{bmatrix}$
$=\begin{bmatrix}1+\frac{a_{\rm{in}}}{a_{\rm{ex}}}&\frac{a_{\rm{in}}}{T_{e}[1]^{\beta}}\\ \frac{T_{e}[1]^{\beta}}{a_{\rm{ex}}}&1\end{bmatrix}\cdots\begin{bmatrix}1+\frac{a_{\rm{in}}}{a_{\rm{ex}}}&\frac{a_{\rm{in}}}{T_{e}[n]^{\beta}}\\ \frac{T_{e}[n]^{\beta}}{a_{\rm{ex}}}&1\end{bmatrix}\begin{bmatrix}T_{e}[n]\\ P[n]\end{bmatrix}$ (11)
where we assume $T_{e}[i]\gg T_{\rm{base}}$ and $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\,(\equiv\beta)$ similar to the assumptions in the main text. Here, we apply the boundary conditions: $P[0]=P_{0}$ and $P[n]=0$. Then, we numerically solve for $T_{e}[i]$ where $i=0,\ldots,n$, given $P_{0}$. Supplementary Fig. 7(b-d) illustrate the case for $n=4$ and $\beta=2$. The results are well fitted by Eq. (5) in the main text. In this figure, we show the results for $\beta=2$, but we have confirmed that it works well for any value of $\beta\geq 0$. These results show that even in a finite and discrete circuit, the parameter dependence can be modeled by Eq. (5) in the main text.
Figure 7: (a) Finite and discrete thermal circuit model, and (b–d) calculation results for the case of $n=4$, where $\beta$ is set to 2, and $a_{\rm{in}}=a_{\rm{ex}}=1$. (b) Linear plots showing the dependence of $T_{e}$ on $P_{0}$, and (c) the corresponding log plots. The colored lines represent the fitting curves based on Eq. (5) in the main text. (d) Best-fit results for $A=ae^{-x/L_{\rm{th}}}$ and $\beta$, where $x=nd$, implying $A=ae^{d/L_{\rm{th}}}e^{-n}\propto e^{-n}$.
Appendix D Validity of CBT setup
We evaluate the validity of the CBT setup. In this measurement, we use the device A with $L_{\rm{cell}}=120$ nm. Supplementary Fig. 8 shows the estimated electron temperature $T_{e}$ obtained by CBT under various temperatures of the mixing chamber $T_{\rm{MXC}}$ in the dilution refrigerator. In the high-temperature regime above 5 K, $T_{\rm{MXC}}$ and $T_{e}$ are expected to be the same. As expected, the results of $T_{\rm{MXC}}=T_{e}$ were obtained, confirming that the temperature estimation using this CBT measurement setup is valid. For the higher temperature regime of 10–-30 K shown in Fig. 5(b) and 5(c) in the main text, the validity is supported by extrapolation from the confirmed data.
Figure 8: Estimated electron temperature $T_{e}$ as a function of the temperature of the mixing chamber $T_{\rm{MXC}}$. Error bar shows the standard deviation error of $T_{e}$ estimate.
We discuss the uncertainty in extracting $T_{e}$ using CBT. Supplementary Fig. 9 illustrates the relationship between the estimated $T_{e}$ and the associated error for three cases, corresponding to Fig. 5(a) and Fig. 5(b) in the main text, and Supplementary Fig. 8, respectively. The error appears to scale proportionally with the estimated $T_{e}$, suggesting that the relative error rate, $\sigma_{T_{e}}/T_{e}$, remains approximately constant throughout this measurement.
Figure 9: Standard deviation error of $T_{e}$ estimate $\sigma_{T_{e}}$ as a function of estimated $T_{e}$ according to (a) Fig. 5(b) and (b) Fig. 5(c) in the main text, and (c) Supplementary Fig. 8, respectevely.